# Guided simulation of conditioned chemical reaction networks

### v2

@@ 1.1. Chemical reaction networks
Let ${\mathbb{S}}\subseteq{\mathbb{Z}}_{\geq 0}^{d}$ denote the state space of the reaction system. A vector in ${\mathbb{S}}$ is of the form $(x_{k})_{k=1}^{d}$, where $x_{k}$ denotes the species count of type $S_{k}$. We study a Markov process $X=\left(X(t)\right)_{t\geq 0}$ on ${\mathbb{S}}$ that models the evolution of species counts over time. That is, at time $t$, the $k$-th element of $X(t)$ is given by #$\#S_{k}(t)$. At a given time $t$, the time until a reaction of type $\ell$ takes place is given by $\tau_{\ell}$. When reaction ^$\hat{\ell}=\operatornamewithlimits{argmin}_{\ell\in{\mathcal{R}}}\tau_{\ell}$ occurs, the process jumps at time $t+\tau_{\hat{\ell}}$ to $X\left(t+\tau_{\hat{\ell}}\right)=X(t)+\xi_{\hat{\ell}}$. A more detailed description of the stochastic model is given in Section 2. The process $X$ evolving on the chemical reaction network is referred to as the chemical reaction process.
@@ 1.2. Statistical problem
Suppose at fixed times $0=t_{0}<t_{1}<\cdots<t_{n}$, we observe $v_{1},\dots v_{n}$, where $v_{k}=L_{k}X(t_{k})$ with $L_{k}\in{\mathbb{R}}^{m_{k}\times d}$ and $m_{k}\leq d$, $k=1,\dots n$. Not assuming $L$ to be the identity matrix is for example important in applications where the measuring device cannot distinguish two or more species, so that only sums of their counts are observed. We will assume the rows of each $L_{k}$ to be linearly independent. Typically reaction rates are unknown and we wish to infer those from the data. Denote the parameter vector containing those rates by $\theta$. Likelihood-based inference for $\theta$ is hampered by the lack of closed-form expressions for the transition probabilities of $X$. However, if the process were observed continuously over time, the problem would be easier. Therefore, it is natural to employ a data-augmentation scheme where we iteratively sample $X$ on $[0,t_{n}]$ conditional on $v_{1},\dots,v_{n}$ and $\theta$ and then update $\theta$ conditional on $X$. In this paper, we focus on the first step, sampling from $\left(X\mid L_{k}X(t_{k})=v_{k},\,k=1,\dots n\right)$. It is a key objective of this paper to show rigorously how this can be done efficiently. Note that a simple rejection sampling scheme where we discard all paths contradicting the observations is valid but very inefficient in most settings.
@@ 1.3. Approach: conditioning by guiding
Our approach builds on earlier work in Corstanje et al., (2023) for general Markov processes in case of a single observation. Let us highlight the main points. The law of the process $X$, conditioned to be in a given state at fixed times, is obtained through Doob’s $h$-transform. That is, there is a function $h$ that depends on the transition probabilities of $X$ that induces a change of measure such that under a new measure, ${\mathbb{P}}^{h}$, the process is conditioned to hit the observed states at times of observation. Since $h$ is typically unknown, we replace it by a fully tractable function $g$ that itself induces a change of measure to a measure ${\mathbb{P}}^{g}$. Under certain conditions, ${\mathbb{P}}^{h}$ is absolutely continuous with respect to ${\mathbb{P}}^{g}$ and
@@
Here, $F$ is known in closed form, depends on $g$ but does not depend on $h$. Weighted samples of the conditioned process can therefore be obtained by sampling under ${\mathbb{P}}^{g}$. The argument for making the above precise is not too hard in case $g$ is bounded and bounded away from zero. However, some natural choices we discuss and have been proposed in the literature require a more delicate argument.
@@ 1.5. Contribution
The proposed methods fit within the framework of Backward Filtering Forward Guiding (van der Meulen & Schauer, (2020)), drawing strongly on techniques for exponential changes of measure as outlined in Palmowski & Rolski, (2002). This enables to construct a guided process that at any time takes into account all future conditionings.
@@ 2. Chemical reaction processes
where ${\mathcal{F}}_{t}^{X}=\sigma(X_{s},\,s\leq t)$. The jump probabilities specified in (2.1) correspond to a process with jumps $(\xi_{\ell})_{\ell\in{\mathcal{R}}}$ and jump rate functions $(\lambda_{\ell})_{\ell\in{\mathcal{R}}}$. Throughout, we impose the following assumptions on the network.
@@
By Theorem 1.22 in Anderson & Kurtz, (2015), there exists a filtration $({\mathcal{F}}_{t})_{t}$ such that for all such $f$
$D^{f}(t):=f(t,X(t))-\int_{0}^{t}(\partial_{t}+{\mathcal{L}})f(s,X(s))\,\mathrm{d}s$
is an $\{{\mathcal{F}}_{t}\}$-martingale. That is, $X$ is the unique solution to the martingale problem for ${\mathcal{A}}:=\partial_{t}+{\mathcal{L}}$.
@@ 2.1. Distribution of reaction times
With each reaction $\ell\in{\mathcal{R}}$, we associate a reaction time $\tau_{\ell}$. The reaction times satisfy, see e.g. Gillespie, (1992),
@@ 2.2. Chemical master equation and chemical Langevin equation
$\,\mathrm{d}N(t)=b_{\mathrm{CLE}}(t,N(t))\,\mathrm{d}t+\sigma_{\mathrm{CLE}}(t,N(t))\,\mathrm{d}W_{\ell}(t),$ (2.5)
@@ 2.3. Examples
    Transcription: $G\to G+M$ with rate $\kappa_{1}>0$.
@@
    Translation: $M\to M+P$ with rate $\kappa_{2}>0$.
@@
    Degradation of mRNA: $M\to\emptyset$ with rate $d_{M}>0$.
@@
    Degradation of protein: $P\to\emptyset$ with rate $d_{P}>0$.
@@
    $S+E\to SE$ with rate $\kappa_{1}$.
@@
    $SE\to S+E$ with rate $\kappa_{2}$.
@@
    $SE\to P+E$ with rate $\kappa_{3}$.
@@ 3.1. First ideas
We first provide some intuition on how the dynamics of a chemical reaction network change upon conditioning on a future event. Let $\mathcal{E}$ denote some event later than time $t$, for example $\{X_{T}=v\}$. For $t\in[0,T)$, $\Delta>0$ and $x\in{\mathbb{S}}$ and $y=x+\xi_{\ell}$,
@@
To sample the conditioned process $h$ is required which is rarely available in closed form. The general idea behind the construction of what we call a guided chemical reaction process is to replace $h$ by a suitable tractable substitute $g$. This gives a process with rates $\lambda_{t}^{g}$. Discrepancies between the true conditioned process and the guided process can be accounted for by evaluating the likelihood ratio of their measures on path space. In this section we will establish sufficient conditions on $g$ for this approach to be valid.
@@ 3.2. Guiding by a change of measure
We assume $x_{0}$ to be known and that =$\mathcal{E}=\{LX(T)=v\}$ is observed for some known $v\in{\mathbb{R}}^{m}$ and a matrix $L\in{\mathbb{R}}^{m\times d}$ of full row rank with $m\leq d$. In case $m=d$, we assume without loss of generality that $L=I$. Additionally, we assume $\mathbb{P}(LX(T)=v\mid X(0)=x_{0})>0$. The extension to multiple observations will follow in a straightforward way in later sections.
@@
As $({\mathcal{F}}_{t}^{X})_{t}=\sigma(X(s):s\leq t)$ is a sub-$\sigma$-algebra of $({\mathcal{F}}_{t})_{t}$, $E^{g}(t)$ is a martingale with respect to $({\mathcal{F}}_{t}^{X})_{t}$ as well whenever $g\in{\mathcal{G}}$.
Let $g\in{\mathcal{G}}$ and let ${\mathbb{P}}_{t}$ denote the law of the process $X$ restricted to ${\mathcal{F}}_{t}^{X}$. Define the new collection of measures $\{{\mathbb{P}}_{t}^{g}:t\in[0,T)\}$ on $(\Omega,{\mathcal{F}}_{t}^{X})$ by
@@
Some authors refer to this process as the twisted process (see e.g. Del Moral & Penev, (2017)). Upon denoting the transition probabilities of $X$ by $p$, i.e. ${\mathbb{P}}(X_{T}\in A\mid X_{t}=x)=\sum_{y\in A}p(t,x;T,y)$ for $A\subseteq{\mathbb{S}}$, it follows from Example 2.4 of Corstanje et al., (2023) that the conditioned process $\left(X\mid LX(T)=v\right)$ is obtained by taking $g=h$ with
@@
Unfortunately, $h$ is intractable as the transition probabilities $p$ are only known in closed-form in very specific cases.
@@
It suffices to show that ${\mathbb{E}}\left[E^{g}(T)\right]=1$. By Lemma 3.1 of Palmowski & Rolski, (2002), $E^{g}$ is a local martingale. Let $\left\{\sigma_{n}\right\}_{n}$ be a localizing sequence for $E^{g}$. Then, if $(E^{g}(t)\,t\in[0,T])$ were uniformly integrable, then
@@
Since we assume throughout $X$ to be nonexplosive, we have that almost surely the maps $t\mapsto g(t,X(t)+\xi_{\ell})/g(t,X(t)$ are almost surely bounded on $[0,T]$ for all $\ell\in{\mathcal{R}}$. The result now follows upon noting that $\int_{0}^{T}\sum_{\ell\in{\mathcal{R}}}\lambda_{\ell}(s,X(s))\,\mathrm{d}s<\infty$ by Assumption (2.1c). ∎
@@ 4. Choices for $g$
It remains to specify the maps $(t,x)\mapsto g(t,x)$ satisfying the assumption of 3.3. Moreover, to be of any use, the event $\{LX_{T}=v\}$ needs to get positive probability under ${\mathbb{P}}^{g}_{T}$.
@@ 4.2. Choosing $g$ using the transition density of a scaled Brownian motion
Note that we have not yet specified $\sigma$. Depending on its choice, $g$ may be radically different from $h$. Nevertheless, it can be used as a change of measure to condition paths of $X$ on the event $\{LX_{T}=v\}$.
@@ 6. Simulation methods
In this section we discuss methods for simulating the guided process on $[0,T]$. This is a Markov jump process with time-dependent intensity. In case $\|LX_{t}-v\|$ is large for $t$ close to $T$, this intensity may blow up.
@@ 6.1.1. Special case: guided process induced by $g_{\varepsilon}$
We consider the guided process with $\lambda_{\ell}^{g_{\varepsilon}}=\alpha^{g_{\varepsilon}}_{\ell}\lambda_{\ell}$, where $\alpha^{g_{\varepsilon}}_{\ell}$ is defined in Equation 4.8. As the map $t\mapsto\alpha_{\ell}^{g_{\varepsilon}}(t,x)$ is monotone, given $(t,x)$, an upper bound can be found at either $(t,x)$ or $(T,x)$. Unfortunately, the upper bound at $T$ is typically far from sharp, hampering efficiency, especially when $\varepsilon$ is chosen to be small. To resolve this issue, we employ Algorithm 2. In lines 22, $\lambda_{\ell}^{g_{\varepsilon}}(\cdot,x)$ is increasing in time. Thus we can choose $\delta$ such that $t+\delta<T$ and use that, on the interval $[t,t+\delta]$, $\lambda_{\ell}^{g_{\varepsilon}}(\cdot,x)$ is upper bounded by $\lambda_{\ell}^{g_{\varepsilon}}(t+\delta,x)$. Hence, on $[t,t+\delta]$, we apply the thinning property for an inhomogeneous Poisson process with rate $\lambda_{\ell}^{g_{\varepsilon}}(\cdot,x)$ and, if no reaction occurs in this interval, we move $t$ to $t+\delta$. Typically, once $LX_{t}=v$ for $t$ near $T$ the map $s\mapsto g(s,x+\xi_{\ell})/g(s,x)$ will be decreasing for choices of $g$ considered in this article. If at line 2 $LX_{t}\neq v$ and $T-(t+\delta)\leq\varepsilon$ we reject the sampled path for specified small $\varepsilon>0$.

### v3

@@ 1.1. Chemical reaction networks
For example, in (1.1), the species set is $\{A,B,C,D\}$ and $\xi$ is given by $(-1,-1,1,1)$. Let ${\mathbb{S}}\subseteq{\mathbb{Z}}_{\geq 0}^{d}$ denote the state space of the reaction system. A vector in ${\mathbb{S}}$ is of the form $(x_{k})_{k=1}^{d}$, where $x_{k}$ denotes the species count of type $S_{k}$. We study a Markov process $X=\left(X(t)\right)_{t\geq 0}$ on ${\mathbb{S}}$ that models the evolution of species counts over time. That is, at time $t$, the $k$-th element of $X(t)$ is given by #$\#S_{k}(t)$. We assume that the initial state $X(0)=x_{0}$ of the process is known. At a given time $t$, the time until a reaction of type $\ell$ takes place is given by $\tau_{\ell}$. When reaction ^$\hat{\ell}=\operatornamewithlimits{argmin}_{\ell\in{\mathcal{R}}}\tau_{\ell}$ occurs, the process jumps at time $t+\tau_{\hat{\ell}}$ to $X\left(t+\tau_{\hat{\ell}}\right)=X(t)+\xi_{\hat{\ell}}$. The reactions are assumed to occur according to an inhomogeneous Poisson process with intensity function that we refer to as the reaction rate. A more detailed description of the stochastic model is given in Section 2. The process $X$ evolving on the chemical reaction network is referred to as the chemical reaction process.
@@ 1.2. Statistical problem
Suppose at fixed times $0<t_{1}<\cdots<t_{n}$, we observe $v_{1},\dots v_{n}$, where $v_{k}=L_{k}X(t_{k})$ with $L_{k}\in{\mathbb{R}}^{m_{k}\times d}$ and $m_{k}\leq d$, $k=1,\dots n$. Not assuming $L$ to be the identity matrix is for example important in applications where the measuring device cannot distinguish two or more species, so that only sums of their counts are observed. We will assume the rows of each $L_{k}$ to be linearly independent. Typically reaction rates are unknown and we wish to infer those from the data. Suppose the reaction rate depend on a parameter vector $\theta$. Likelihood-based inference for $\theta$ is hampered by the lack of closed-form expressions for the transition probabilities of $X$. However, if the process were observed continuously over time, the problem would be easier. Therefore, it is natural to employ a data-augmentation scheme where we iteratively sample $X$ on $[0,t_{n}]$ conditional on $v_{1},\dots,v_{n}$ and $\theta$ and then update $\theta$ conditional on $X$. In this paper, we focus on the first step, sampling from $\left(X\mid L_{k}X(t_{k})=v_{k},\,k=1,\dots n\right)$. It is a key objective of this paper to show rigorously how this can be done efficiently. Note that a simple rejection sampling scheme where we discard all paths contradicting the observations is valid but very inefficient in most settings.
@@ 1.3. Approach: conditioning by guiding
Our approach builds on earlier work in Corstanje et al., (2023) for general Markov processes in case of a single observation. Let us highlight the main points. The law of the process $X$, conditioned to be in a given state at fixed times, is obtained through Doob’s $h$-transform. That is, there is a function $h\colon[0,T]\times{\mathbb{S}}\to\mathbb{R}_{+}$ that depends on the transition probabilities of $X$ which induces a change of measure. Under the new measure, ${\mathbb{P}}^{h}$, the process is conditioned to hit the observed states at times of observation. Since $h$ is typically unknown, we replace it by a fully tractable function $g\colon[0,T]\times{\mathbb{S}}\to\mathbb{R}_{+}$ that itself induces a change of measure to a measure ${\mathbb{P}}^{g}$. Under certain conditions, ${\mathbb{P}}^{h}$ is absolutely continuous with respect to ${\mathbb{P}}^{g}$ and
@@
Here, $F$ is known in closed form, depends on $g$ but does not depend on $h$. Weighted samples of the conditioned process can therefore be obtained by sampling under ${\mathbb{P}}^{g}$. The argument for making the above precise is not too hard in the case where $g$ is bounded and bounded away from zero. However, some natural choices we discuss and have been proposed in the literature require a more delicate argument.
@@ 1.5. Contribution
The proposed methods fit within the framework of Backward Filtering Forward Guiding (van der Meulen & Schauer, (2020)), drawing strongly on techniques for exponential changes of measure as outlined in Palmowski & Rolski, (2002). This enables us to construct a guided process that at any time takes into account all future conditionings.
@@ 2. Chemical reaction processes
where ${\mathcal{F}}_{t}^{X}=\sigma(X^{t})$. The jump probabilities specified in (2.1) correspond to a process with jumps $(\xi_{\ell})_{\ell\in{\mathcal{R}}}$ and jump rate functions $(\lambda_{\ell})_{\ell\in{\mathcal{R}}}$. Throughout, we impose the following assumptions on the network.
@@
By Theorem 1.22 in Anderson & Kurtz, (2015), there exists a filtration $({\mathcal{F}}_{t})_{t\geq 0}$ such that for all such $f$
$D^{f}(t):=f(t,X(t))-\int_{0}^{t}(\partial_{s}+{\mathcal{L}})f(s,X(s))\,\mathrm{d}s$
is an $({\mathcal{F}}_{t})$-martingale. That is, $X$ is the unique solution to the martingale problem for ${\mathcal{A}}:=\partial_{t}+{\mathcal{L}}$.
@@ 2.1. Distribution of reaction times
To the $\ell$-th reaction, we associate a reaction time $\tau_{\ell}$, with distribution specified by
@@ 2.2. Chemical master equation and chemical Langevin equation
$\,\mathrm{d}Y(t)=b_{\mathrm{CLE}}(t,Y(t))\,\mathrm{d}t+\sigma_{\mathrm{CLE}}(t,Y(t))\,\mathrm{d}W_{\ell}(t),\qquad Y(0)=x_{0}$ (2.5)
@@ 2.3. Examples
    Transcription: $G\to G+M$ with rate constant $\kappa_{1}$.
@@
    Translation: $M\to M+P$ with rate constant $\kappa_{2}>0$.
@@
    Degradation of mRNA: $M\to\emptyset$ with rate constant $d_{M}>0$.
@@
    Degradation of protein: $P\to\emptyset$ with rate constant $d_{P}>0$.
@@
    $S+E\to SE$ with rate constant $\kappa_{1}$.
@@
    $SE\to S+E$ with rate constant $\kappa_{2}$.
@@
    $SE\to P+E$ with rate constant $\kappa_{3}$.
@@ 3.1. First ideas
We first provide some intuition on how the dynamics of a chemical reaction network change upon conditioning on a future event. Let $\mathcal{E}$ denote some event later than time $t$, for example $\{X(T)=v\}$. For $t\in[0,T)$, $\Delta>0$ and $x\in{\mathbb{S}}$ and $y=x+\xi_{\ell}$,
@@
To sample the conditioned process, the function $h$ is required which is rarely available in closed form. The general idea behind the construction of what we call a guided chemical reaction process is to replace $h$ by a suitable tractable substitute $g$. This gives a process with rates $\lambda_{t}^{g}$. Discrepancies between the true conditioned process and the guided process can be accounted for by evaluating the likelihood ratio of their measures on path space. In this section we will establish sufficient conditions on $g$ for this approach to be valid.
@@ 3.2. Guiding by a change of measure
We assume $x_{0}$ to be known and that =$\mathcal{E}=\{LX(T)=v\}$ is observed for some known $v\in{\mathbb{R}}^{m}$ and a matrix $L\in{\mathbb{R}}^{m\times d}$ of full row rank with $m\leq d$. In case $m=d$, we assume without loss of generality that $L=I$. Additionally, we assume $\mathbb{P}(LX(T)=v\mid X(0)=x_{0})>0$. The extension to multiple observations will follow in a straightforward way in later sections. Let ${\mathcal{A}}$ denote the infinitesimal generator of the space-time process $(t,X(t))$. That is $({\mathcal{A}}g)(s,x)=\lim_{\Delta\downarrow 0}\Delta^{-1}\mathbb{E}\left[g(s+\Delta,X(s+\Delta))-g(s,X(s))\mid X(s)=x\right]$, for those $g$ (that map to $\mathbb{R}$) for which the limit exists. Such $g$ are said to be in the domain of the generator, denoted by ${\mathcal{D}}({\mathcal{A}})$. While ${\mathcal{D}}({\mathcal{A}})$ is implicitly defined, as part of the definition of ${\mathcal{A}}$, we can find a more descriptive description on a smaller set of functions. For those functions ${\mathcal{A}}=\partial_{t}+{\mathcal{L}}$, where ${\mathcal{L}}$ is defined in (2.3). From this, one can see that for instance functions $g$ that are differentiable with respect to their time-argument will typically be in the domain.
@@
As $({\mathcal{F}}_{t}^{X})_{t}=\sigma(X(s):s\leq t)$ is a sub-$\sigma$-algebra of $({\mathcal{F}}_{t})_{t}$, $E^{g}(t)$ is a martingale with respect to $({\mathcal{F}}_{t}^{X})_{t}$ as well whenever $g\in{\mathcal{G}}$. In 3.3, we give an explicit condition under which $g\in{\mathcal{G}}$.
Let $g\in{\mathcal{G}}$ and let ${\mathbb{P}}_{t}$ denote the law of the process $X$ restricted to ${\mathcal{F}}_{t}^{X}$. We define a new collection of measures $\{{\mathbb{P}}_{t}^{g}:t\in[0,T)\}$ by defining ${\mathbb{P}}_{t}$ on $(\Omega,{\mathcal{F}}_{t}^{X})$ by
@@
Some authors refer to this process as the twisted process (see e.g. Del Moral & Penev, (2017)). Upon denoting the transition probabilities of $X$ by $p$, i.e. ${\mathbb{P}}(X(T)\in A\mid X(t)=x)=\sum_{y\in A}p(t,x;T,y)$ for $A\subseteq{\mathbb{S}}$, it follows from Example 2.4 of Corstanje et al., (2023) that the conditioned process $\left(X\mid LX(T)=v\right)$ is obtained by taking $g=h$ with
@@
Unfortunately, $h$ is intractable as the transition probabilities $p$ are only known in closed-form in very specific cases including 2.2.
@@
It suffices to show that ${\mathbb{E}}\left[E^{g}(T)\right]=1$. By Lemma 3.1 of Palmowski & Rolski, (2002), $E^{g}$ is a local martingale. Let $\left\{\sigma_{n}\right\}_{n}$ be a localizing sequence for $E^{g}$. Then, if $(E^{g}(t),\,t\in[0,T])$ were uniformly integrable, then
@@
Since we assume throughout $X$ to be nonexplosive, we have that almost surely the maps $t\mapsto g(t,X(t)+\xi_{\ell})/g(t,X(t))$ are almost surely bounded on $[0,T]$ for all $\ell\in{\mathcal{R}}$. The result now follows upon noting that $\int_{0}^{T}\sum_{\ell\in{\mathcal{R}}}\lambda_{\ell}(s,X(s))\,\mathrm{d}s<\infty$ by Assumption (2.1c). ∎
@@ 4. Choices for $g$
It remains to specify the maps $(t,x)\mapsto g(t,x)$ satisfying the assumption of 3.3. Moreover, to be of any use, the event $\{LX(T)=v\}$ needs to get positive probability under ${\mathbb{P}}^{g}_{T}$.
@@ 4.2. Choosing $g$ using the transition density of a scaled Brownian motion
Note that we have not yet specified $\sigma$. Depending on its choice, $g$ may be radically different from $h$. Nevertheless, it can be used as a change of measure to condition paths of $X$ on the event $\{LX(T)=v\}$.
@@ 6. Simulation methods
In this section we discuss methods for simulating the guided process on $[0,T]$. This is a Markov jump process with time-dependent intensity. In case $\|LX(t)-v\|$ is large for $t$ close to $T$, this intensity may blow up.
@@ 6.1.1. Special case: guided process induced by $g_{\varepsilon}$
We consider the guided process with $\lambda_{\ell}^{g_{\varepsilon}}=\alpha^{g_{\varepsilon}}_{\ell}\lambda_{\ell}$, where $\alpha^{g_{\varepsilon}}_{\ell}$ is defined in Equation 4.8. As the map $t\mapsto\alpha_{\ell}^{g_{\varepsilon}}(t,x)$ is monotone, given $(t,x)$, an upper bound can be found at either $(t,x)$ or $(T,x)$. Unfortunately, the upper bound at $T$ is typically far from sharp, hampering efficiency, especially when $\varepsilon$ is chosen to be small. To resolve this issue, we employ Algorithm 2. In lines 22, $\lambda_{\ell}^{g_{\varepsilon}}(\cdot,x)$ is increasing in time. Thus we can choose $\delta$ such that $t+\delta<T$ and use that, on the interval $[t,t+\delta]$, $\lambda_{\ell}^{g_{\varepsilon}}(\cdot,x)$ is upper bounded by $\lambda_{\ell}^{g_{\varepsilon}}(t+\delta,x)$. Hence, on $[t,t+\delta]$, we apply the thinning property for an inhomogeneous Poisson process with rate $\lambda_{\ell}^{g_{\varepsilon}}(\cdot,x)$ and, if no reaction occurs in this interval, we move $t$ to $t+\delta$. Typically, once $LX(t)=v$ for $t$ near $T$ the map $s\mapsto g(s,x+\xi_{\ell})/g(s,x)$ will be decreasing for choices of $g$ considered in this article. If at line 2 $LX(t)\neq v$ and $T-(t+\delta)\leq\varepsilon$ we reject the sampled path for specified small $\varepsilon>0$.